The risk factor and time – Compounding and discounting Suppose you have just celebrated your 19th...

  1. The risk factor and time – Compounding and discounting

Suppose you have just celebrated your 19th birthday. A rich uncle set up a trust fund for you that will pay you $100,000 when you turn 25. If the relevant discount rate is 11 percent, how is this fund worth today?

Answer: ______________ (explain your answer

You have been offered an investment that will double your money in 12 years. What rate of return are you being offered? Check your answer using the rule of 72.

Answer:
Answer #1

Hi as per policy we will solve only first question here.

1) Here Future Value F = $100,000

time t = 25-19 = 6

discount rate r = 11%

Fund worth today = Present Value P=?

P = F/(1+r)^t

P = 100000/(1+11%)^6

P = 100000/1.11^6

P = $53,464.08

Thanks